3. Find the area of the shaded region enclosed by the following functions y = 2 - 1x1 y = -1
Answer #1

Shade area = g-11)-y -) --- c) 9 Q Lejryci, Jose 3 B-1) 뼛 > x=(1) dx 게게 왕가 =오3가-22 - 4 (:( |-x Sac [] 한 시에 g9 = 338) 3 _313